40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3218     Entry time: Wed Jul 14 12:31:11 2010
Author: nancy 
Type: Update 
Category: General 
Subject: Weekly Update 

Summary of this week's work Wednesday - Aligned the mode cleaner with Koji, and then measured the beam characteristics at MC2 end. Koji then taught me how to read the WFS signals Thursday - wrote a script to measure the signals and calculated the coefficients relating mirror movement and DC signals of WFS. To know the possibility of the control, found SVD of the coeff matrix, and condition number. Friday - Set up the measurement of QPD linear response using a laser outside the cavity. Took data. Monday - did the calculations and plotting for the above experiment. Then played around with the MEDM screens , and also tried to see what happens to the Power Spectrum of WFS signals by changing the coefficients in teh matrix. (failed) Tuesday - played around with WFS, tried seeing what it does when switched on at different points, and also what it does when I disturb the system while WFS has kept it locked.

Today - had switched off the WFS sensors yesterday night after locking the MC as wanted to know that how does MC behave when no WFS gain is applied. I checked in the morning, the MC was locked all night. I am now proceding with my calculations for the sensing matrix
ELOG V3.1.3-